### v1

# Chi-Hsi Kung

Abstract
Understanding a procedural activity requires modeling both how action steps transform the scene, and how evolving scene transformations can influence the sequence of action steps, even those that are accidental or erroneous. Yet, existing work on procedure-aware video representations fails to explicitly learned the state changes (scene transformations). In this work, we study procedure-aware video representation learning by incorporating state-change descriptions generated by LLMs as supervision signals for video encoders. Moreover, we generate state-change counterfactuals that simulate hypothesized failure outcomes, allowing models to learn by imagining the unseen “What if” scenarios. This counterfactual reasoning facilitates the model’s ability to understand the cause and effect of each step in an activity. To verify the procedure awareness of our model, we conduct extensive experiments on procedure-aware tasks, including temporal action segmentation, error detection, and more. Our results demonstrate the effectiveness of the proposed state-change descriptions and their counterfactuals, and achieve significant improvements on multiple tasks. Full paper [11] is available at https://arxiv.org/abs/2503.21055.
Figure 1: Illustration of our learning framework. (a) We train frame features by temporally contrasting neighboring and distant frames by incorporating Before, After, and State-change Counterfactuals (SC-CF). (b) We align clip features with narration features. (c) Clip features are aggregated with the aggregator (Agg) as video features. Then the video features are contrasted positively with summaries and negatively with Missing-step Counterfactuals (K-CF) and Misordered Counterfactuals (M-CF). Note that all text features are extracted with the frozen text encoder.
[…]
1 Introduction
Understanding procedure activities from video data is essential for a variety of applications, including video retrieval [26], intelligent collaborative agents [7], and robot learning from human demonstration [21]. In contrast to general video action recognition that focuses on a single step in a short clip [25, 10], procedure-aware video understanding requires capturing both the “what changed” (actual action-induced state transformations) [23], such as Before and After state shown in Figure 1 (a), and “what could have changed” (hypothetical deviations) [30], such as State-change counterfactuals shown in Figure 1 (a) and Missing-step and Misordered shown in Figure 1 (c). These capabilities are key to understanding long-form procedures with sequentially dependent steps.
There have been many approaches proposed to learn procedure-aware representations, including learning spatiotemporal features [4, 25], using action labels as supervision [27, 31], incorporating temporal order of steps [31], and consulting external activity procedure knowledge databases [32]. However, these methods often fail to model how scene states evolve or could have evolved under different action outcomes.
We propose a hierarchical video representation learning framework for procedure understanding that leverages state changes and counterfactuals generated by an LLM [3]. At the clip level, we model before, after, and counterfactual states to capture local action-induced changes. We use temporal contrastive learning to bring after states close to later frames’ features and push apart before and counterfactuals. At the video level, we extend this with long-form counterfactuals like missing-step and misordered, improving procedure awareness.
We evaluate the learned procedure-aware video representations in three key procedural video understanding tasks—error detection, temporal action segmentation, and action phase classification and frame retrieval—and show that they significantly enhance performance compared to strong baselines. Our contributions are summarized as follows:
  1. 1.
    A video representation method leveraging state changes and counterfactuals for procedural understanding.
  2. 2.
    A hierarchical framework aligning frame-, clip-, and video-level features using these descriptions.
  3. 3.
    Our method achieves state-of-the-art results on procedure-aware tasks with detailed analysis.
2 Related Work
2.1 Video Procedure-Aware Representations
Recently, several papers have attempted to align video features in datasets such as HowTo100M [17] or Ego4D [6] with text descriptions extracted through ASR [18, 16], refined subtitles [15] with the external database WikiHow [9, 16], or manually labeled annotations [14, 31, 1]. However, these approaches fail to explicitly learn action-induced state changes or detect deviations like erroneous steps, leading to overfitting on correctly executed actions and limiting procedural understanding. We address this by incorporating state-change descriptions and counterfactuals to model action transformations and hypothetical failures for improved procedure awareness.
2.2 State-Change Understanding
Visual state changes have been widely studied [13, 22, 19], and we draw inspiration from SCHEMA [19] in using LLMs to generate before- and after-state descriptions. Our work differs in three key ways: (1) we target general video representation learning beyond a single task; (2) we incorporate counterfactuals for causal reasoning; and (3) our descriptions capture state changes of objects, humans, and environments—not just interacted objects.
3 Pretraining Objective:
State Change & Counterfactual
Here, we present our pretraining strategy that incorporates the generated state changes and state-change counterfactuals to learn procedure-aware representations. We build upon HierVL [1], a framework that learns hierarchical video-language representations at two temporal scales, clip-level and video-level. Please refer to [1] for an overview of the framework. We extend HierVL with finer-grained frame-level alignment.
Frame-Level Alignment At the frame level, the model is supervised by our proposed Before state loss $\mathcal{L}_{\text{before}}$ and After-state loss $\mathcal{L}_{\text{after}}$, enhancing action-induced transformation for procedural understanding. The mathematical formulation of each is [8],
$\mathcal{L}$ $=\frac{1}{|B|}\sum_{i\in B}\frac{-1}{|P(i)|}\sum_{p\in P(i)}\log\frac{\exp(f_{i}^{T}z_{p}/\tau)}{\sum_{n\in N(i)}\exp(f_{i}^{T}z_{n}/\tau)},$ (1)
where $B$ is the batch size, $f_{i}$ is the visual embedding of the $i^{\text{th}}$ frame, $z_{j}$ is either a visual or text embedding, $\tau$ is a temperature hyperparameter, and $P(i)$ and $N(i)$ denote the positive and negative samples of the $i^{\text{th}}$ frame, respectively. Given a set of $K=4$ frames sub-sampled from a video-clip, $L_{\text{before}}$ aims to align earlier-in-time frames along with the before-state text embeddings, while pushing them apart from later-in-time frames, the after-state, and counterfactual text embeddings. On the other hand, $L_{\text{after}}$ aims to align later-in-time frames to the after state while separating them from earlier-in-time frames, the before-state, and counterfactual text embeddings.
Clip-Level Alignment At the clip-level, we leverage the $L_{v2t}$ loss described in [1, 14], which seeks to align the video-clip embeddings to their corresponding text narrations from EgoClip [14]. Note that since we do not train a text encoder, the symmetric $L_{t2v}$ loss is neglected here. For more details on this loss, please see [1, 14]. The resulting loss for the first two scales is therefore
$\mathcal{L}_{\text{child}}=\mathcal{L}_{v2t}+\lambda(\mathcal{L}_{\text{before}}+\mathcal{L}_{\text{after}}),$ (2)
where $\lambda$ is a hyperparameter controlling the strength of the state-change aware supervision.
Video-Level Alignment The goal of this loss is to align video-level visual embeddings to summary text embeddings and to enhance procedural awareness using video-level counterfactuals. We first obtain video-level visual embeddings from clip-level embeddings by leveraging a self-attention block as an aggregator function. Then, using contrastive learning, each visual embedding is aligned to its corresponding summary text embedding and contrasted against text embeddings of the generated counterfactual Misordered and Missing-step. The formulation is defined as,
$\mathcal{L}_{\text{parent}}$ $=-\sum_{i\in B}\log\frac{\sum_{p\in P(i)}\exp(V_{i}^{T}S_{p})}{\sum_{n\in N(i)}\exp(V_{i}^{T}S_{n})+\exp(V_{i}^{T}S_{n,w}^{cf})},$ (3)
where $V_{i}$ is the aggregated video-level visual embedding, $S_{j}$ is a summary text embedding and $S_{j,w}^{cf}$ are Misordered and Missing-step counterfactual text embeddings where $w\in\{1,\ldots,W\}$ is the total number of counterfactuals used, and $P(i)$ and $N(i)$ denote the positive and negative samples of the $i^{\text{th}}$ video, respectively. We perform video-level alignment as in Eq. (3) after every 5 mini-batches of clip- and frame-level alignment as in Eq. (2).
4 Experiments
Tasks We evaluate our model’s learned representations on four procedural video-understanding tasks: temporal action segmentation on GTEA [5] and EgoPER [12], error detection on EgoPER [12], and action phase classification and frame retrieval on the Align-Ego-Exo dataset [28]. We report F1 score, mean average precision (mAP), edit distance, frame accuracy, or Error Detection Accuracy (EDA) metrics based on each task convention.
Implementation Details We use ASFormer [29], EgoPED [12], an SVM, and nearest-neighbors for each task, respectively, and equip them with different input representation features. In the Align-Ego-Exo dataset, we merge the validation and test sets for more robust results.
Baselines We compare to I3D and CLIP features [2, 20] commonly used in long-form video tasks and procedure-aware representations with publicly available pretrained model weights, including MIL-NCE [18], and PVRL [31]. In addition, we evaluate the VLM HierVL [1] with only its video encoder.
4.1 Experimental Results
Temporal Action Segmentation In Table 1, we find that MIL-NCE performs significantly worse than others, which we hypothesize is due to the poor quality of description made by ASR. Furthermore, we compare our model with the VLM HierVL. Even though HierVL is trained with both video and text encoders, which require considerably more computational resources, our model outperforms HierVL in most metrics on both datasets. The performance gap against all non-VLM models is even larger.
Error Detection In Table 2, we observe that all procedure-aware representations outperform general visual representations, such as I3D and CLIP, highlighting the importance of procedure awareness in the context of error detection. Our proposed method surpasses the state-of-the-art in the EDA metric [12], even outperforming the VLM HierVL.
Action Phase Classification & Frame Retrieval In Table 3 and Table 4, our learned representations outperform other models on average across actions in both phase classification and retrieval, on both the ego+exo views (regular) and egocentric only (ego) settings, while also achieving superior or competitive performance on most actions individually.
Table 1: Temporal action segmentation results on the GTEA [5] and EgoPER [12] datasets.
GTEA EgoPER
Method F1@10 F1@25 F1@50 Edit Acc F1@50 Edit Acc
I3D [2] 90.1 88.8 79.2 84.6 79.7 48.8 71.9 73.9
CLIP [20] 88.5 86.2 77.6 87.1 75.6 44.2 71.2 70.8
MIL-NCE [18] 67.9 61.3 44.6 67.9 58.3 47.3 69.1 73.6
PVRL [31] 85.2 82.6 72.2 81.1 71.2 45.6 73.2 73.4
HierVL [1] 90.4 88.5 81.2 86.7 78.5 52.6 73.0 77.3
Ours 89.8 89.1 81.6 86.8 80.0 54.4 74.1 79.0
Table 2: Error detection results on EgoPER [12]. HTM denotes the HowTo100M [17] dataset. Text denotes the VLM model with a trainable text encoder. The metric is EDA.
Method Pretraining Data Quesadilla Oatmeal Pinwheel Coffee Tea All
Random - 19.9 11.8 15.7 8.20 17.0 14.5
I3D [2] Kinetics 62.7 51.4 59.6 55.3 56.0 57.0
CLIP [20] WIT [24]+Text 77.6 69.6 66.9 68.5 75.6 71.6
MIL-NCE [18] HTM 77.3 69.8 65.7 68.0 69.8 70.1
PVRL [31] HTM 75.7 71.2 65.5 67.5 76.4 71.3
HierVL [1] Ego4D+Text 77.9 70.8 65.2 67.4 75.1 71.3
Ours Ego4D 78.9 71.6 68.3 68.3 76.6 72.7
Table 3: Action phase classification results on the Align-Ego-Exo dataset [28]. The metric is F1 score.
Break Eggs Pour Milk Pour Liquid Tennis Forehand All
Method regular ego regular ego regular ego regular ego regular ego
CLIP [20] 50.1 54.9 50.4 49.8 61.3 63.7 76.3 78.2 59.5 61.6
MIL-NCE [18] 45.5 45.0 45.9 44.2 61.2 65.3 59.5 62.3 53.0 54.2
PVRL [31] 54.6 60.6 51.6 46.6 63.0 69.0 68.2 74.5 59.4 62.7
Ours 56.2 65.8 48.1 47.6 68.1 70.6 72.7 75.1 61.3 64.8
Table 4: Frame retrieval results on the Align-Ego-Exo dataset [28]. The metric is mAP@10.
Break Eggs Pour Milk Pour Liquid Tennis Forehand All
Method regular ego regular ego regular ego regular ego regular ego
CLIP [20] 63.5 68.0 59.3 59.2 55.9 56.1 79.1 88.7 64.4 68.0
MIL-NCE [18] 58.0 57.4 47.3 51.0 57.7 59.2 74.8 84.3 59.5 63.0
PVRL [31] 59.5 63.1 58.3 59.3 50.2 55.1 78.3 88.9 61.6 66.6
Ours 66.5 69.4 51.4 54.9 62.4 67.8 79.4 88.9 64.9 70.3
5 Conclusions
We present a novel procedure-aware video representation learning framework that first incorporates state-change descriptions and state-change counterfactuals in clip-level alignment, enhancing causal reasoning of action transformations. Then, it utilizes video-level counterfactuals that perturb the local actions and create hypothesized scenarios to facilitate the understanding of activity procedures. Our learned representations demonstrate strong effectiveness in terms of procedure awareness and achieve state-of-the-art results on several benchmarks.